# A Genealogy of Foundation Models in Remote Sensing

### v2

@@ 9. Combined Approaches
Contrastive Radar-Optical Masked Auto-encoders (CROMA) (Fuller et al., 2023) is one such model. CROMA trains on Sentinel-1 and Sentinel-2 imagery and trains an encoder for each sensor. However, CROMA takes inspiration from the Fast Language Image-Pretraining (FLIP) framework (Jia et al., 2021), which independently masks each mode of data before feeding them into encoders and performing contrastive learning in the standard CLIP (Radford et al., 2021) setup. Masking images before encoding them results in a significantly less memory intensive process than encoding the raw image when vision transformers are used. Therefore, the FLIP framework enables contrastive training with a much larger batch size with the same or less memory footprint. This results in faster convergence. CROMA implements this approach with remotely sensed images, contrastively learning between masked Sentinel-1 and Sentinel-2 images. CROMA also introduces a third encoder that takes in concatenated Sentinel-1 and Sentinel-2 encodings to generate multi-sensor encodings. These fused encodings are fed into a multi-sensor decoder responsible for constructing the initial Sentinel-1 and Sentinel-2 images. In this manner, CROMA is able to introduce contrastive and reconstruction loss terms into their framework. This results in a model that effectively learns uni-sensor representations for Sentinel-1 and Sentinel-2 and multi-sensor representations of both Sentinel-1 and Sentinel-2. CROMA’s framework is benchmarked against BigEarthNet (Sumbul et al., 2019) and FMoW (Christie et al., 2018) classification tasks, demonstrating SOTA performance with images from a single sensor and multi-sensed images. BigEarthNet contains roughly half a million Sentinel-2 images where each patch can have multiple land-cover classes. CROMA also utilizes modified Attention with Linear Biases (ALiBi) positional embeddings (Press et al., 2022) for their ViT encoders. These embeddings provide an advantage over standard sinusoidal positional embeddings by biasing the embeddings based on Euclidean distance between key-query pairs, thus imbuing the embeddings with a degree of spatial awareness.
@@
Figure 9. An illustration of shared and unique concepts among RS foundation models which utilize hybrid learning strategies that combine contrastive, MIM, and distillation approaches.:
@@ 10.2. Under-Tapped Modes for Geospatial Representation Learning
Much of the current research focuses on drawing parallels to the work done on CV foundation models. Given the success of foundation models in the CV domain and the sheer amount of imagery that is available, be it from satellites or other aerial sources, this is a necessary first step. However, models such as MMCPP (Chen et al., 2023) use the CLIP contrastive learning objective and rely on other modalities, such as trajectory maps, points of interest, and aerial images. Their success should serve as a somewhat cautionary tale against becoming too invested in foundation models that purely focus on satellite and/or aerial imagery. Environmental stations measure features such as groundwater, weather, or snow melt data. They provide important climate and atmospheric information at a very high temporal resolution, often recording data on a daily or even hourly basis. More human-centric data, such as social media posts, social connectedness indices (Bailey et al., 2018), and trajectory data, also provide invaluable information for understanding more societally-focused downstream tasks. Both forms of information are not observable from space, but could provide useful modes due to their different perspective and high temporal resolution. Data on rapidly changing phenomena also tends to not be observable from space. Changes can occur too quickly for a satellite to capture due to orbit and revisit time. GAIR’s (Liu et al., 2025) usage of INR representations for satellite imagery also provides a useful starting point for mapping granular, ground-level data to satellite imagery.
@@ 10.3. Performance versus Training Expenses
Aside from a carbon footprint perspective, models with huge numbers of parameters have their downstream applications somewhat limited due to computational requirements. Not all users have the hardware necessary to fine-tune foundation models with a high parameter count. Some models, such as MOSAIKS (Rolf et al., 2021), focus on lightweight frameworks in order to promote access to remote sensing foundation models. MOSAIKS has competitive results on a number of downstream tasks, indicating that increasing parameters within RS foundation models may follow a law of diminishing returns. On the other hand, the encoder-decoder structure that is predominant in foundation models allows for repeated, diverse usage of representations on downstream tasks. Arguably, leveraging significant resources to generate high quality embeddings as a one-time cost is the most effective path moving forward, as this approach results in SOTA performance on downstream tasks with relatively few resources necessary for future fine-tuning.

### v3

@@ 9. Combined Approaches
Contrastive Radar-Optical Masked Auto-encoders (CROMA) (Fuller et al., 2023) is one such model. CROMA trains on Sentinel-1 and Sentinel-2 imagery and trains an encoder for each sensor. However, CROMA takes inspiration from the Fast Language Image-Pretraining (FLIP) framework (Jia et al., 2021), which independently masks each mode of data before feeding them into encoders and performing contrastive learning in the standard CLIP (Radford et al., 2021) setup. Masking images before encoding them results in a significantly less memory intensive process than encoding the raw image when vision transformers are used. Therefore, the FLIP framework enables contrastive training with a much larger batch size with the same or less memory footprint. This results in faster convergence. CROMA implements this approach with remotely sensed images, contrastively learning between masked Sentinel-1 and Sentinel-2 images. CROMA also introduces a third encoder that takes in concatenated Sentinel-1 and Sentinel-2 encodings to generate multi-sensor encodings. These fused encodings are fed into a multi-sensor decoder responsible for constructing the initial Sentinel-1 and Sentinel-2 images. In this manner, CROMA is able to introduce contrastive and reconstruction loss terms into their framework. This results in a model that effectively learns uni-sensor representations for Sentinel-1 and Sentinel-2 and multi-sensor representations of both Sentinel-1 and Sentinel-2. CROMA’s framework is benchmarked against BigEarthNet (Sumbul et al., 2019) and FMoW (Christie et al., 2018) classification tasks, demonstrating SOTA performance with images from a single sensor and multi-sensed images at the time of publication in 2023. BigEarthNet contains roughly half a million Sentinel-2 images where each patch can have multiple land-cover classes. CROMA also utilizes modified Attention with Linear Biases (ALiBi) positional embeddings (Press et al., 2022) for their ViT encoders. These embeddings provide an advantage over standard sinusoidal positional embeddings by biasing the embeddings based on Euclidean distance between key-query pairs, thus imbuing the embeddings with a degree of spatial awareness.
@@
Figure 9. An illustration of shared and unique concepts among RS foundation models which utilize hybrid learning strategies that combine contrastive, MIM, and distillation approaches.
A sort of Venn diagram where a variety of RS models that utilize hybrid learning approaches are compared with one another to highlight similarities and differences.
@@
Dynamic One-For-All (DOFA) (Xiong et al., 2024) builds on GFM’s approach of distilling against a CV model pre-trained on ImageNet while also creating a framework that can take in data from any number of different sensors. DOFA follows a hypernetwork approach (Chauhan et al., 2024), where a hypernetwork takes in the wavelength details of a given band and then generates weights that are used to initialize a neural network. Image bands are masked, then fed into that neural network to generate encodings. The encodings are then fed through an encoder-decoder structure, with a mirrored hypernetwork at the end responsible for reconstructing the original image. This framework is designed to learn bandwidth specific features via the hypernetwork and shared features via the shared encoder-decoder. DOFA is trained on SAR imagery from Sentinel-1, multispectral imagery from Sentinel-2, hyperspectral imagery from enMAP, and high resolution RGB imagery from NAIP. For the distillation objective, DOFA generates embeddings from the RGB channels (if present) and compares those embeddings to those of an ImageNet pre-trained teacher model on the same RGB bands. In the case of SAR imagery, a random polarization is selected for the third channel to ensure the correct input shape for the CV teacher model. Similar to GFM, the authors of DOFA cite a desire for reduced C02 emissions and efficiency during pre-training (Xiong et al., 2024) as a motivation for the distillation objective. The resulting framework achieves SOTA performance on downstream tasks such as BigEarthNet (Sumbul et al., 2019), So2Sat (Zhu et al., 2019), and EuroSat (Helber et al., 2019) at the time of publication in 2024.
@@ 10.2. Under-Tapped Modes for Geospatial Representation Learning
Many current RS foundation models build on CV foundation model frameworks. Given the success of foundation models in the CV domain and the sheer amount of imagery that is available, be it from satellites or other aerial sources, this is a necessary first step. However, some models have achieved success by training heavily on non-image modes of data. For example, MMCPP (Chen et al., 2023) uses the CLIP contrastive learning objective and trains on population trajectory maps, points of interest, and aerial RGB images. Such models make a compelling case to avoid purely focusing on satellite and/or aerial imagery when training RS foundation models.
Environmental stations measure features such as groundwater, weather, or snow melt data. They provide important climate and atmospheric information at a very high temporal resolution, often recording data on a daily or even hourly basis. More human-centric data, such as social media posts, social connectedness indices (Bailey et al., 2018), and trajectory data, also provide invaluable information for understanding more societally-focused downstream tasks. Both forms of information are not observable from space, but could provide useful modes due to their different perspective and high temporal resolution. Data on rapidly changing phenomena also tends to not be observable from space. Changes can occur too quickly for a satellite to capture due to orbit and revisit time. GAIR’s (Liu et al., 2025) usage of INR representations for satellite imagery also provides a useful starting point for mapping granular, ground-level data to satellite imagery.
@@ 10.3. Performance versus Training Expenses
Aside from a carbon footprint perspective, models with huge numbers of parameters have their downstream applications somewhat limited due to computational requirements. Not all users have the hardware necessary to fine-tune foundation models with a high parameter count. Some models, such as MOSAIKS (Rolf et al., 2021), focus on lightweight frameworks in order to promote access to remote sensing foundation models. MOSAIKS has competitive results on a number of downstream tasks such as forest cover prediction(Hansen et al., 2013) at the time of publication in 2021, indicating that increasing parameters within RS foundation models may follow a law of diminishing returns. On the other hand, the encoder-decoder structure that is predominant in foundation models allows for repeated, diverse usage of representations on downstream tasks. Arguably, leveraging significant resources to generate high quality embeddings as a one-time cost is the most effective path moving forward, as this approach results in SOTA performance on downstream tasks with relatively few resources necessary for future fine-tuning.